The sum of emily and mikayla's ages is 12 and the diffrence of their ages is 6. if emily is older what are their ages
Answer by TimothyLamb(4379)

x = emily age
y = mikayla age
---
x + y = 12
x - y = 6
---
put the system of linear equations into standard form
---
x + y = 12
x - y = 6
---
copy and paste the above standard form linear equations in to this solver:
http://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 9
y = 3
---
Free algebra tutoring live chat:
http://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
http://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
http://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
http://sooeet.com/math/system-of-linear-equations-solver.php
---